OpenStudy (anonymous):

Perform indicated operation 1. 2a/3 divided by 1/2 2. 5/7-1/2 Please help

OpenStudy (anonymous):

\[\huge \cfrac{\cfrac{2a}{3}}{\cfrac{1}{2}}=\frac{2\cdot 2a}{3\cdot 1}\] go on...

OpenStudy (anonymous):

When you divide by a fraction you take the reciprocal and multiply, so 2/1 *2a/3 For the second one you have to make sure they have the same denominators, then subtract.

OpenStudy (anonymous):

\[\LARGE \frac57-\frac12=\frac{5\cdot 2}{7\cdot 2}-\frac{?}{7\cdot 2}\] ....

OpenStudy (anonymous):

Would I multiply the 5x2=10, would I put that over the 7.2

OpenStudy (anonymous):

it is not 7.2 is 7*2 or \[7\times 2\]

OpenStudy (anonymous):

sorry thats what i meant

OpenStudy (anonymous):

\[\LARGE \frac{10}{14}-\frac{7\cdot 1}{14}= ?\]

OpenStudy (anonymous):

3/14

OpenStudy (anonymous):

that's correct. Well done ! ... now tell me about the first one....what do you get ?

OpenStudy (anonymous):

Im sorry i dont get it

OpenStudy (anonymous):

\[\LARGE \cfrac{2\cdot 2a}{3\cdot 1}=\frac{4a}{3}=\frac43\cdot a\]

OpenStudy (anonymous):

you can't simplify it any further ...

OpenStudy (anonymous):

thanks for all your help